Female: Mean: 592 Std. dev.: 94 Male: Mean:628 Std. dev.:89 Calculate the interval corresponding to one,...

Female:
Mean: 592
Std. dev.: 94


Male:
Mean:628
Std. dev.:89

Calculate the interval corresponding to one, two, and three standard deviations.  Type your work showing how you obtained these intervals. Keep the endpoints of the final intervals as whole numbers and clearly label and list these three intervals in your document as shown below:
68% interval (lower value, upper value)
           95% interval (lower value, upper value)
           99.7% interval (lower value, upper value)
Answer #1

Female:

68% Interval = (592 - 94, 592 + 94) = (498, 686)

95% Interval = (592 - 2*94, 592 + 2*94) = (404, 780)

99.7% Interval = (592 - 3*94, 592 + 3*94) = (310, 874)

Male:

68% Interval = (628 - 89, 628 + 89) = (539, 717)

95% Interval = (628 - 2*89, 628 + 2*89) = (450, 806)

99.7% Interval = (628 - 3*89, 628 + 3*89) = (361, 895)